### v1

# Smooth hybrid inflation in light of ACT DR6 data

Abstract
Smooth hybrid inflation is a hybrid inflation model which is free from topological defects and predicts the density perturbation with the spectral index of about $0.97$. We show that the prediction on the spectral index is robust regardless the power of nonrenormalizable terms and meets with the latest results reported by Atacama Cosmology Telescope. Viable scenarios for baryogenesis and dark matter are also discussed.
preprint: EPHOU-25-011
I Introduction
The inflationary cosmology has been the standard paradigm of modern cosmology. It solves various problems in the standard big bang cosmology Starobinsky:1979ty ; Sato:1981qmu ; Guth:1980zm ; Linde:1981mu ; Albrecht:1982wi and explains the origin of the density perturbation in our Universe from the quantum nature of an inflaton field Hawking:1982cz ; Starobinsky:1982ee ; Guth:1982ec . The scalar perturbation is generated from the inflaton’s quantum fluctuation during quasi de Sitter inflationary expansion. As a result, the spectrum is almost scale invariant, in fact its scale dependence, parameterized by the spectral index, reflects the dynamics of inflation. The scalar spectral index has been reported as $n_{s}\simeq 0.965$ from the measurement by Planck Planck:2018vyg . Thus, it has been regarded that so called Starobinsky inflation Starobinsky:1980te , Higgs inflation Bezrukov:2007ep and $\alpha$-attractor inflation Kallosh:2013hoa ; Ferrara:2013rsa ; Kallosh:2013yoa in the class of pole inflation Galante:2014ifa ; Broy:2015qna are promising models of inflation, because those models predict $n_{s}\simeq 0.96$.
Recently, the latest results from Atacama Cosmology Telescope (ACT) ACT:2025fju ; ACT:2025tim have been released and the scalar spectral index of $n_{s}\simeq 0.974$ has been reported ACT:2025fju for after combining data of ACT, Planck and baryon acoustic oscillation. This shift of estimated $n_{s}$ motivates us to reconsider models for inflation  Kallosh:2025rni ; Aoki:2025wld ; Dioguardi:2025vci ; Gialamas:2025kef ; Salvio:2025izr ; Antoniadis:2025pfa ; Dioguardi:2025mpp ; Rehman:2025fja ; Gao:2025onc ; He:2025bli ; Drees:2025ngb ; Yin:2025rrs ; McDonald:2025odl ; Byrnes:2025kit ; Maity:2025czp ; Yi:2025dms ; Gialamas:2025ofz ; Yogesh:2025wak ; Haque:2025uga ; Pallis:2025nrv ; Wolf:2025ecy ; McDonald:2025tfp . Although hybrid inflation models Linde:1993cn ; Copeland:1994vg can be constructed within realistic particle physics models Lyth:1998xn , in particular, in supersymmetric model Dvali:1994ms , those have appeared not to be promising, because those generally predict too large spectral index. In addition, hybrid inflation models suffer from the formation of topological defects, because the spontaneous symmetry breaking by the waterfall field terminates inflation with generating topological defects. The typical mass per unit length of cosmic strings formed after hybrid inflation confronts with the present observational bound on it Battye:2010xz ; Planck:2013mgr .
In this article, we revisit smooth hybrid inflation which has been proposed as a hybrid inflation model free from topological defects and predicts spectral index of about $0.97$ Lazarides:1995vr . We show that the prediction on the spectral index is robust regardless the power of nonrenormalizable terms and well meets with the latest ACT DR6 data.
II Smooth hybrid inflation
The superpotential for a smooth hybrid inflation is given by
$W=S\left(-\mu^{2}+\frac{(\bar{\Psi}\Psi)^{m}}{M^{2m-2}}\right),$ (1)
where this model is based on certain gauge group $G$ under which $S$ is a singlet, and $\Psi$ and $\bar{\Psi}$ are a pair of a vector-like representation. We may consider $U(1)_{B-L}$ as an example. While $m=2$ is assumed in the original paper Lazarides:1995vr , we generalize that $m$ is an integer ($\geq 2$) as in Refs. Yamaguchi:2004tn ; Kawasaki:2006zv . With the minimal Kahler potential, the scalar potential in supergravity in the Planck unit is
$V=$ $e^{|S|^{2}+|\Psi|^{2}+|\bar{\Psi}|^{2}}\left[\left|-\mu^{2}+\frac{(\bar{\Psi}\Psi)^{m}}{M^{2m-2}}\right|^{2}(1-|S|^{2}+|S|^{4})\right.$
$+|S|^{2}\left|m\bar{\Psi}\frac{(\bar{\Psi}\Psi)^{m-1}}{M^{2m-2}}+\Psi^{*}\left(-\mu^{2}+\frac{(\bar{\Psi}\Psi)^{m}}{M^{2m-2}}\right)\right|^{2}$
$\left.+|S|^{2}\left|m\Psi\frac{(\bar{\Psi}\Psi)^{m-1}}{M^{2m-2}}+\bar{\Psi}^{*}\left(-\mu^{2}+\frac{(\bar{\Psi}\Psi)^{m}}{M^{2m-2}}\right)\right|^{2}\right]+D{\rm-terms}$
$\simeq$ $\left(-\mu^{2}+\frac{\psi^{2m}}{M^{2m-2}}\right)^{2}\left(1+\frac{|S|^{4}}{2}+2\psi^{2}\right)+2|S|^{2}\psi^{2}\left(m\frac{\psi^{2m-2}}{M^{2m-2}}-\mu^{2}\right)^{2}.$ (2)
Here, in the last expression, the D-flat condition ($\Psi=\bar{\Psi}^{*}\equiv\psi$) are imposed, and the third terms in the F-term of $\Psi$ and $\bar{\Psi}$ are neglected compared to the first terms, and the lowest order supergravity corrections for $S$ and $\psi$ only are kept in the F-term of $S$, since we are interested in the field region $|S|^{2},\psi^{2}\ll 1$. The supersymmetric global minimum is located at
$\left(S,\psi\right)=\left(0,\mu^{\frac{1}{m}}M^{\frac{m-1}{m}}\right),$ (3)
where the mass of inflatons is
$m_{S}^{2}=m_{\psi}^{2}\simeq 2m^{2}\mu^{2}\left(\frac{\mu}{M}\right)^{\frac{2(m-1)}{m}}.$ (4)
The stationary condition for $\psi$ is
$V,_{\psi}\simeq 4\psi\left(-\mu^{2}+m\frac{\psi^{2m-2}}{M^{2m-2}}\right)\left(-\mu^{2}+\frac{\psi^{2m}}{M^{2m-2}}+m(2m-1)|S|^{2}\frac{\psi^{2m-2}}{M^{2m-2}}\right),$ (5)
and we find a nontrivial local minimum
$|S|^{2}\left(\frac{\psi}{M}\right)^{2m-2}=\frac{\mu^{2}}{m(2m-1)},$ (6)
for a large $|S|$ as
$|S|^{2}>\frac{\psi^{2}}{m(2m-1)}.$ (7)
In this region, inflationary expansion can be realized due to the false vacuum energy $V\simeq\mu^{4}$. For this minimum for $\psi$, we can see
$V_{,\psi\psi}\simeq\frac{8\mu^{2}}{(2m-1)|S|^{2}}\left[(m-1)\mu^{2}+mM^{2}\left(\frac{\mu^{2}}{m(2m-1)|S|^{2}}\right)^{\frac{m}{m-1}}\right]\gg\mu^{4}.$ (8)
for $|S|\ll 1$. Hence, we consider that the $\psi$ field can trace the instantaneous minimum Eq. (6). By substituting Eq. (6) into Eq. (2), we obtain the potential for only the canonical inflaton $\sigma=|S|/\sqrt{2}$ as
$V\simeq\mu^{4}\left(1-\frac{2M^{2}}{\mu^{2}}\frac{m-1}{2m-1}\left(\frac{2\mu^{2}}{m(2m-1)}\right)^{\frac{m}{m-1}}\sigma^{-\frac{2m}{m-1}}+\frac{\sigma^{4}}{8}\right).$ (9)
Yamaguchi and Yokoyama verified this procedure Yamaguchi:2005qm . The final term in the potential comes from a supergravity effect. The slow roll parameters are defined by
ϵ:=assignitalic-$\epsilon:=$ $\frac{1}{2}\left(\frac{V_{,\sigma}}{V}\right)^{2},$ (10)
$\eta:=$ $\frac{V_{,\sigma\sigma}}{V},$ (11)
where the derivatives of the potential are given as
$\frac{V_{,\sigma}}{V}\simeq\frac{2M^{2}}{\mu^{2}}\frac{2m}{2m-1}\left(\frac{2\mu^{2}}{m(2m-1)}\right)^{\frac{m}{m-1}}\sigma^{-\frac{2m}{m-1}-1}+\frac{\sigma^{3}}{2},$ (12)
$\frac{V_{,\sigma\sigma}}{V}\simeq-\frac{2M^{2}}{\mu^{2}}\frac{2m}{2m-1}\frac{3m-1}{m-1}\left(\frac{2\mu^{2}}{m(2m-1)}\right)^{\frac{m}{m-1}}\sigma^{-\frac{2m}{m-1}-2}+\frac{3}{2}\sigma^{2}.$ (13)
We define the transition point $\sigma_{t}$ in the potential from Eq. (12) as
$\sigma_{t}^{\frac{2m}{m-1}+4}=\frac{4M^{2}}{\mu^{2}}\frac{2m}{2m-1}\left(\frac{2\mu^{2}}{m(2m-1)}\right)^{\frac{m}{m-1}}.$ (14)
While the potential slope is dominantly given by the $\sigma^{-\frac{2m}{m-1}}$ term for $\sigma<\sigma_{t}$, the supergravity effect is significant for $\sigma>\sigma_{t}$. The critical point where the slow roll condition breaks down as $\eta(\sigma_{c})=-1$ is given by
$\sigma_{c}^{\frac{2m}{m-1}+2}=\frac{2M^{2}}{\mu^{2}}\frac{2m}{2m-1}\frac{3m-1}{m-1}\left(\frac{2\mu^{2}}{m(2m-1)}\right)^{\frac{m}{m-1}}.$ (15)
The number of e-folds during inflation is expressed as
$N=\frac{3m-1}{4m-2}\left(\frac{\sigma}{\sigma_{c}}\right)^{2\frac{2m-1}{m-1}},$ (16)
as long as
$\sigma(N)<\sigma_{t}$ (17)
is satisfied. This constrains the parameters of the model as shown in Fig. 1 where each contours are of $\sigma(N)/\sigma_{t}$ for $N=60$ and the gray shaded region is not compatible with the condition (17). If the condition (17) violates, supergravity effects are not negligible and the resultant $n_{s}$ exceeds $1$ Panagiotakopoulos:1997qd ; Linde:1997sj hence not compatible with observation.
Figure 1: The contours of $\sigma(N)/\sigma_{t}$ for $N=60$. The gray shaded region violates the condition (17).
The amplitude of the density perturbation is estimated as
${\cal P}_{\zeta}=\frac{\mu^{4}}{12\pi^{2}}\left(\frac{m-1}{3m-1}\sigma_{c}\right)^{-2}\left(\frac{4m-2}{3m-1}N\right)^{\frac{3m-1}{2m-1}},$ (18)
and the spectral index is expressed as
$n_{s}=1-\frac{3m-1}{2m-1}\frac{1}{N}.$ (19)
One can immediately realize that, for $m\geq 2$, the spectral index is more red tilted than that of the standard supersymmetric hybrid inflation with a logarithmic correction in the potential. The tensor to scalar ratio is given by
$r=16\epsilon,$ (20)
with
ϵ=italic-$\epsilon=$ $\frac{1}{2}\left(\frac{m-1}{3m-1}\left(\frac{\sigma_{c}}{\sigma}\right)^{\frac{3m-1}{m-1}}\sigma_{c}\right)^{2}.$ (21)
In addition, the mass of the inflaton in the vacuum is given by
$m_{\sigma}^{2}=2M^{2}\left(\frac{\mu}{M}\right)^{2/m}\left[-\mu^{2}+m\left(\frac{\mu}{M}\right)^{2-2/m}\right]^{2}.$ (22)
II.1 $m=2$ case
We note some analytic formulas for $m=2$. The condition (17) for a supergravity effect to be negligible is rewritten as
$1\gg(4\sqrt{3}N^{2}\mu M)^{2/3}=\left(\frac{N}{55}\right)^{4/3}\left(\frac{\mu M}{4.8\times 10^{-5}}\right)^{2/3}.$ (23)
The amplitude of the density perturbation reads
${\cal P}_{\zeta}=20\times 10^{-10}\left(\frac{\mu^{5}/M}{10^{-33/2}}\right)^{2/3}\left(\frac{N}{55}\right)^{5/3},$ (24)
and the spectral index is expressed as
$n_{s}=1-\frac{5}{3}\frac{1}{N}\simeq 0.97\quad{\rm for}\quad N\simeq 55.$ (25)
Then, from Eqs. (23) and (24), we find that the scale of $\mu$ must be somewhat smaller than a typical GUT scale as $\mu<3.4\times 10^{-4}\simeq 8.2\times 10^{14}$ GeV. The tensor to scalar ratio is smaller than $\mathcal{O}(10^{-7})$ as shown in Fig. 2.
Figure 2: The cut-off scale $M$ dependence on the tensor to scalar ratio for $m=2$. The thicker line is for $N=60$, while the thinner line is for $N=50$.
II.2 General $m$ cases
We show how results change if we vary $m$ in the range of $2\leq m\leq\infty$. The results for fixed $M=10^{-3}$ are shown in Fig. 3 of the $n_{s}-r$ plane. The thicker (thinner) blue curve is the prediction for $N=60$ $(50)$ as $m$ varies from $2$ to $\infty$, where $\mu$ is adjusted to normalize the amplitude of density perturbation given in Eq. (18). The orangeish and purplish shadings stands for the likelihood contours for Planck-LB-BK18 and P-ACT-LB-BK18, respectively, reported in Ref. ACT:2025tim .
Figure 3: The model prediction for $M=10^{-3}$. The thicker blue curve is for $N=60$, while the thinner blue curve is for $N=50$.
II.3 Post-inflationary evolution
After inflation, the inflaton fields oscillate around the minimum (3). The inflaton decay may be caused through the superpotential $W=ySH_{u}H_{d}$ where $y$ is a coupling constant, and $H_{u}$ and $H_{d}$ are superfields of the up-type and down-type Higgs doublet fields, respectively. The reheating temperature after inflation $T_{R}$ is controlled by the magnitude of $y$. If $\Psi$ and $\bar{\Psi}$ are charged under $U(1)_{B-L}$, $\Psi$ or $\bar{\Psi}$ can couple with right-handed neutrino superfield $N$, for instance, as
$W=y_{N}\Psi NN.$ (26)
If $T_{R}$ is large enough, the baryon asymmetry could be generated by thermal leptogenesis Fukugita:1986hr ; Davidson:2008bu . On the other hand, if $T_{R}$ is not so high, the baryon asymmetry can be generated by Affleck-Dine mechanism Affleck:1984fy . If the superpotential (26) exists, nonthermal leptogenesis by the inflaton decay is also possible Asaka:1999yd . In fact, Jeannerot et al. studied nonthermal leptogenesis in the original smooth hybrid inflation Jeannerot:2001qu . As is suggested by the Higgs boson mass of $125$ GeV, soft SUSY breaking mass parameters seem to be large Okada:1990vk ; Okada:1990gg ; Ellis:1990nz ; Haber:1990aw ; Ellis:1991zd . Thus, the promising candidate of dark matter in this scenario would be Higgsino with the mass of about $1$ TeV left due to thermal freeze out Cirelli:2007xd .
III Conclusion
We revisited smooth hybrid inflation model whose spectral index of the density perturbation is about $0.97$, which well agrees with the latest results reported by ACT. By treating the power of non-renormalizable terms as a free parameter, we have found that the predicted $n_{s}$ lies within the uncertainty of the latest ACT results for any value of its power $m$. The scenario seems to be compatible with various baryogenesis mechanism and Higgsino-like neutralino dark matter.
Acknowledgments
This work is supported in part by the U.S. DOE Grant No. DE-SC0012447 and DE- SC0023713 (N.O.) and KAKENHI Grants No. JP23K03402 (O.S.).